### v1

# Multi Source COVID-19 Detection via Kernel-Density-based Slice Sampling

Abstract
We present our solution for the Multi-Source COVID-19 Detection Challenge, which classifies chest CT scans from four distinct medical centers. To address multi-source variability, we employ the Spatial-Slice Feature Learning (SSFL) framework with Kernel-Density-based Slice Sampling (KDS). Our preprocessing pipeline combines lung region extraction, quality control, and adaptive slice sampling to select eight representative slices per scan. We compare EfficientNet and Swin Transformer architectures on the validation set. The EfficientNet model achieves an F1-score of 94.68, compared to the Swin Transformer’s 93.34. The results show the performance of our KDS-based pipeline on multi-source data and highlight the importance of dataset balance in multi-institutional medical imaging evaluation.
1 Introduction
The COVID-19 pandemic has increased interest in automated diagnostic tools to support clinical decision-making. Chest computed tomography (CT) can reveal pulmonary manifestations such as ground-glass opacities and bilateral infiltrates associated with COVID-19 [6, 7].
Developing automated COVID-19 detection systems from CT scans faces challenges in multi-institutional settings. CT data from different hospitals exhibit variability due to differences in acquisition parameters, scanner models, imaging protocols, and patient demographics. This domain shift can affect deep learning model performance when deployed across different medical centers [9, 8, 10].
To address these challenges, we participate in the Multi-Source COVID-19 Detection Challenge, which requires binary classification of COVID/Non-COVID cases using CT scans from four diverse medical institutions. The heterogeneous nature of this dataset presents unique opportunities to develop and validate domain-robust algorithms that can generalize across different clinical environments.
Our approach builds upon the SSFL framework [4, 3, 2, 5], incorporating Kernel-Density-based Slice Sampling (KDS) to handle multi-source variability effectively. The key insight is that while CT scans from different institutions may vary in acquisition parameters and quality, the underlying anatomical structures and pathological patterns remain consistent. By focusing on diagnostically relevant lung regions and intelligently sampling representative slices, our method can learn domain-invariant features while maintaining computational efficiency.
Through experiments comparing EfficientNet [12] and Swin Transformer [11] architectures, we evaluate the effectiveness of our preprocessing pipeline on multi-source data. Our results contribute to understanding the trade-offs between model complexity and preprocessing design in multi-institutional medical imaging tasks.
2 Methodology
Figure 1: Pipeline overview of the proposed multi-source COVID-19 detection approach, showing the preprocessing steps (SSFL and KDS) and model architectures.
We present our solution for the Multi-Source COVID-19 Detection Challenge, which addresses the classification of chest CT scans across data from four distinct medical centers. Our approach builds upon the SSFL framework [4, 3, 2, 5], incorporating preprocessing and sampling techniques to handle multi-source variability. Figure 1 illustrates the overall architecture of our approach.
2.1 Spatial-slice Feature Learning
To address the heterogeneity across different medical centers, we implement a preprocessing pipeline for data consistency.
Lung Region Extraction: We first apply minimum filtering with a $3\times 3$ kernel to the original CT images to smooth noise and suppress artifacts. Following this, fixed-threshold binarization with threshold $t$ and hole filling are performed to separate lung regions from the background. This step identifies diagnostically relevant content across different scanning protocols.
The filtered image $Z_{\text{filtered}}$ is obtained through:
$Z_{\text{filtered}}(i,j)=\frac{\sum_{p=-k}^{k}\sum_{q=-k}^{k}w(p,q)\times Z(i+p,j+q)}{\sum_{p=-k}^{k}\sum_{q=-k}^{k}w(p,q)}$ (1)
where $w(p,q)$ represents the weight at position $(p,q)$ in the filter kernel with $k=1$. Afterwards, we perform consistency checks on image dimensions within each CT scan folder. Scans with inconsistent image sizes or containing fewer than 5 slices are excluded to maintain data integrity. This quality control step ensures that only complete scans are used for subsequent analysis.
Lung Region Cropping: Based on the identified lung regions, we determine cropping boundaries using threshold-based segmentation with empirically assigned threshold $t$:
$\text{Mask}[i,j]=\begin{cases}0,&\text{if }Z_{\text{filter}}[i,j]<t\\ 1,&\text{if }Z_{\text{filter}}[i,j]\geq t\end{cases}$ (2)
This process removes background areas and focuses on lung tissue regions across different medical centers. All cropped images are resized to $256\times 256$ pixels for model input. Figure 2 shows an example of lung region cropping, demonstrating the focus on relevant areas after preprocessing.
2.2 Kernel-Density-based Slice Sampling
Following the identification of CT slice ranges, we apply our KDS method to select representative slices. This approach addresses varying scan lengths and slice quality across different sources.
Distribution Estimation: We apply Kernel Density Estimation (KDE) to estimate the distribution of lung areas within the slice range:
$\hat{f}_{h}(x)=\frac{1}{sh}\sum_{i=1}^{s}K\left(\frac{x-x_{i}}{h}\right)$ (3)
where $h$ is the bandwidth constant calculated by Scott’s rule with $h=1.06\cdot\sigma\cdot n^{-1/5}$, $s=100$ for smoothing factor, and $K$ is a Gaussian kernel:
$K(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{2\sigma^{2}}\right)$ (4)
where $\sigma^{2}=h^{2}/2$ for the kernel bandwidth.
Adaptive Sampling: The Cumulative Distribution Function (CDF) is calculated and partitioned into 8 percentile intervals:
[…]
We sample exactly 8 representative slices per scan, with the number of slices from each interval proportional to its contribution to the overall CDF. This provides coverage of the lung anatomy while maintaining computational efficiency. Figure 3 shows examples of CT slices selected through our KDS method.
Figure 2: Example of a CT slice before and after lung region cropping. Left: original slice. Right: cropped slice.
Figure 3: Example of representative CT slices selected using KDS.
3 Experiment Results
3.1 Dataset Description
The dataset used in this study was obtained from the multi-source COVID-19 Detection Challenge. These cases originate from four distinct source data centers, labeled numerically from 0 to 3. All CT scans were categorized into COVID-19 and non-COVID-19 classes.
The training set consists of 564 COVID-19 scans and 659 non-COVID-19 scans. The validation set included 128 COVID-19 scans and 180 non-COVID-19 scans. The test set contained 1,488 cases, for which the class labels were not provided and needed to be predicted. The overall dataset statistics are summarized in Table 1.
Table 1: Distribution of CT scan and slice counts across different data sources and COVID-19 status for training, validation, and testing sets.
(a) Scan-level counts
Type Source COVID non-COVID Total
Training 0 175 164 339
1 175 165 340
2 39 165 204
3 175 165 340
Validation 0 43 45 88
1 43 45 88
2 0 45 45
3 42 45 87
Total 0 218 209 427
1 218 210 428
2 39 210 249
3 217 210 427
Testing 1,488
(b) Slice-level counts
Type Source COVID non-COVID Total
Training 0 48,740 45,989 94,729
1 8,143 10,988 19,131
2 10,771 47,846 58,617
3 30,834 26,974 57,808
Validation 0 10,957 11,909 22,866
1 2,070 2,814 4,884
2 0 13,283 13,283
3 7,739 8,942 16,681
Total 0 59,697 57,898 117,595
1 10,213 13,802 24,015
2 10,771 61,129 71,900
3 38,573 35,916 74,489
Testing 306,203
3.2 Experimental Setup and Details
All experiments were conducted using the PyTorch framework on NVIDIA RTX 3090 GPUs. Implementation details will be introduced as follows:
Model Architecture: We compare two architectures in our experiments: EfficientNet-B7 [12, 4] and Swin Transformer-Base [11]. The EfficientNet model is based on the EfficientNet-B7 backbone with a classification head, while the Swin Transformer uses the base configuration with hierarchical feature learning. Both models process the 8 selected representative slices from each CT scan to make final predictions through global average pooling and fully connected classification layers.
Training Configuration: We employ EfficientNet-B7 as our backbone architecture. The model is trained using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and weight decay of $5\times 10^{-4}$. The training batch size is set to 20, validation batch size to 1, and training is performed for up to 3 epochs with early stopping based on validation performance. We use binary cross-entropy as the loss function with mixed precision training (AMP) for computational efficiency. All input images are resized to $256\times 256$ pixels and normalized using ImageNet statistics.
Data Augmentation: To improve generalization and robustness, we apply several data augmentation techniques during training:
  • Horizontal flipping with 50% probability
  • Shift-scale-rotate transformation with shift limit 0.2, scale limit 0.2, and rotation within ±30 degrees (50% probability)
  • Hue, saturation, and value adjustment with shift limits of ±0.2 (50% probability)
  • Random brightness and contrast adjustment within ±0.2 range (50% probability)
  • Coarse dropout with 20% probability for regularization
  • Normalization using ImageNet statistics [1]
For validation, we apply only resizing to $256\times 256$ pixels and normalization to ensure consistent evaluation conditions.
Cross-Validation Strategy: We apply stratified 5-fold cross-validation for the training set. This approach maintains the class distribution of COVID-19 and non-COVID-19 cases within each fold, preserving the proportion of positive and negative samples as in the original dataset. In each cross-validation round, the training folds were combined with the validation set, resulting in a total of five folds per round.
Evaluation Metrics: We use F1-score as our primary evaluation metric, which provides a balanced measure of precision and recall. Additionally, we report AUC-ROC scores to assess the model’s performance across different operating points.
3.3 Experiment Results
We evaluate the performance of both EfficientNet-B7 and Swin Transformer models under two different training strategies: source-specific training and multi-source joint training.
Source-Specific Training: Table 2 presents the performance when each model is trained separately on individual sources and evaluated on the corresponding validation sets. The EfficientNet-B7 model achieves F1-scores of 97.53 and 90.90 on Sources 0 and 3, respectively. The Swin Transformer achieves F1-scores of 93.17 and 94.15 on Sources 0 and 3. Both models show lower performance on Source 2, where the validation set contains only positive samples.
Multi-Source Joint Training: When training on all sources combined, as shown in Table 3, the EfficientNet-B7 model achieves an overall F1-score of 94.68, compared to the Swin Transformer’s 93.34. The EfficientNet-B7 model achieves an AUC-ROC of 0.9813 compared to 0.9797 for the Swin Transformer.
The lower performance on Source 2 for both models can be attributed to the validation set containing only positive samples (COVID cases), which affects the calculation of binary classification metrics. This indicates the importance of balanced validation sets for performance evaluation in multi-source medical imaging tasks.
Table 2: Performance comparison across different hospitals between the proposed EfficentNet-B7 and Swin Transformer on the COVID-19-CT-DB validation set.
Model type Hospital AUC ROC F1-score
Swin Transformer 0 0.9922 93.17
1 0.9515 87.17
2 NA 100
3 0.9698 94.15
EfficentNet-B7 0 0.9981 97.53
1 0.9564 87.96
2 NA 49.21
3 0.9672 90.90
Results affected by class imbalance (validation set contains only covid samples).
Table 3: Overall performance comparison between the proposed EfficentNet-B7 and Swin Transformer on the COVID-19-CT-DB validation set (multi-source joint training).
Model type AUC ROC F1-score
Swin Transformer 0.9797 93.34
EfficentNet-B7 0.9813 94.68
Our results show several findings: (1) The EfficientNet-B7 model outperforms the Swin Transformer in our evaluation scenarios, (2) Both models show performance variation across sources, with lower scores on Source 2, and (3) The class imbalance in Source 2’s validation set affects performance evaluation, highlighting considerations for dataset balance in multi-institutional studies.
4 Conclusion
We present a robust solution for the Multi-Source COVID-19 Detection Challenge using the SSFL framework with KDS strategy. Our EfficientNet model achieves an F1-score of 94.68, outperforming the Swin Transformer’s 93.34, while reducing data redundancy by 70%. The results demonstrate that well-designed 2D architectures with effective preprocessing can achieve superior performance compared to complex transformer-based approaches, providing a practical foundation for AI-assisted COVID-19 detection in clinical settings.